# Vulnerabilities of smart contracts and mitigation schemes: A Comprehensive Survey

### v1

@@ Abstract
Ethereum smart contracts are highly powerful; they are immutable and retain massive amounts of tokens. However, smart contracts keep attracting attackers to benefit from smart contract flaws and Ethereum’s unexpected behaviour. Thus, methodologies and tools have been proposed to help implementing secure smart contracts and to evaluate the security of smart contracts already deployed. Most related surveys focus on tools without discussing the logic behind them; in addition, they assess the tools based on papers rather than testing the tools and collecting community feedback. Other surveys lack guidelines on how to use tools specific to smart contract functionalities.
This paper presents a literature review combined with an experimental report, that aims to assist developers in developing secure smarts, with a novel emphasis on the challenges and vulnerabilities introduced by NFT fractionalization by addressing the unique risks of dividing NFT ownership into tradeable units called fractions. It provides a list of frequent vulnerabilities and corresponding mitigation solutions. In addition, it evaluates the community’s most widely used tools by executing and testing them on sample smart contracts. Finally, a complete guidance on how to secure smart contracts is presented.
@@
Blockchain, Ethereum, Smart contracts, Formal verification, Semantic verification, Fuzzing, Software security, Software quality.=-15pt
I Introduction
@@
Blockchain technology has exploded in popularity over the years due to its immutability, security, and transparency in a permissionless and decentralized environment. One of the well-known and used implementations is Ethereum. As of 31 May 2023, the Ethereum cryptocurrency’s market capital surpassed $220 billion, with millions of transactions being executed everyday [1]. Ethereum uses Turing Complete blockchain technology [2], allowing developers to implement smart contracts. Smart contracts are executable programs stored in the blockchain written mostly in solidity [3]. Using smart contracts facilitates the execution of pre-defined terms without consulting third parties in an anonymous, transparent and tempered residence way.
@@
The Ethereum blockchain is the foundation for numerous financial projects. Examples include: (a) platforms for decentralised finance (DeFi), which enable users to access a variety of financial services like lending and borrowing, trading, and insurance without the need for intermediates. (b) Stablecoins, cryptocrencies whose value is tied to that of a fiat currency or other asset, like gold or the US dollar, (c) Security token offerings (STOs), which are virtual assets that simulate ownership of a physical asset like a stock or piece of real estate. (d) Non-fungible tokens (NFTs), which are unique digital assets that might signify ownership of a physical or digital object, like a work of art or a collectible, (e) Amidst this variety, NFT fractionalization emerges as a noteworthy trend, dividing ownership of NFTs into smaller, more accessible units, thus broadening participation in the digital asset market. This value attracts attackers to exploit different vulnerabilities related to the implementation of smart contracts in order to steal cryptocurrencies or tamper with assets. For instance, a recent attack was on May 2023 where Level Finance Exchange announced the loss of more than 214,000 $LVL tokens, approximated to be $1.01 million [4], caused by an attack on its smart contract by manipulating a recursive calling vulnerability. There is a variety of vulnerabilities and causes such as arithmetic overflows, reentrancy, inadequate randomness, calling an unknown third-party contract code and failing to check the return status of external calls. The runtime environment that supports contract code execution provides more attack points. One example is when malevolent miners pick and choose which transactions are included in a mined block, or in what order they are included.
@@
Since vulnerabilities in smart contracts can result in significant financial loss, multiple tools are proposed to identify vulnerabilities in Solidity smart-contracts. Detection methods include: (a) Symbolic execution when the program is abstractly executed to cover multiple possible inputs. (b) Fuzz testing by injecting several invalid inputs. (c)Taint analysis by checking an input flow. (d) Formal verification when the behavior of the smart contract is checked mathematically using a formal model.
@@
In this paper we conduct a literature review combined with an experimental report. We present the most common vulnerabilities in Solidity smart contracts, including those posed by the fractionalization of NFTs and the corresponding mitigation schemes,. We also overview and compare the most popular schemes and tools used to detect vulnerabilities of smart contracts. Finally, we propose a set of guidelines for auditing smart contracts.
@@
In this survey, we searched 2 popular databases Engineering Village and Scopus, in addition to web articles written by the Ethereum community Concerning vulnerabilities in Solidity smart-contracts, mitigation tools and guidelines. Moreover, we experimentally studied 5 tools widely used to detect vulnerabilities in smart contracts namely Oyente [5], Slither [6], Mythril [7], Manticore [8] and Echidna [9].
@@
The rest of this paper is organised as follows: Section II outlines the methodology we used to produce this survey. Section III presents the most common vulnerabilities of smart contracts, and the corresponding mitigation schemes. Section IV presents common methodologies used to detect smart contract vulnerabilities. Section V introduces the most useful tools to detect vulnerabilities in smart contracts. Section VI presents a set of guidelines for auditing smart contracts. Section VII presents an overview of existing related surveys. Finally, Section VIII concludes the paper.
II Survey Methodology
@@
RQ3: What are the existing methodologies used to detect the vulnerabilities of smart contracts and how do they compare with each other?
@@
To address these questions, we conducted a Multivocal Literature Review (MLR)[10] for data preparation; This technique incorporates both gray literature and white literature. Gray literature includes blogs, videos, and forums; it is usually written by practitioners from industry and academia. It’s not peer reviewed. White literature includes peer-reviewed research articles from journals. We carried out this step by choosing the published studies in journals, conferences with high impact factor and competitive acceptance rate. We also check the citation count of the studies being chosen on Google Scholar to evaluate their impact on the evolution of this emerging paradigm. The rationale behind choosing MLR is that the field of smart contracts security is still relatively new; thus, including literature from practitioners is helpful to have a complete overview. Moreover, the feedback provided by smart contract developers is crucial in selecting the most useful tools.
@@
We defined a search string presented in Listing 1 to collect the following resources: (a) Review articles about smart contract security; (b) Papers that discuss vulnerabilities in Ethereum smart contracts; (c) Paper that that cover detection methodologies or mitigation schemes of smart contract vulnerabilities; and (d) Papers about tools that detect vulnerabilities in smart contacts. We have only included paper that contains the keywords in either the subject, title or abstract. The term ”opcode” refers to the basic instructions carried out by the blockchain’s virtual machine, which regulate the logic and operations of the smart contract. Turing completeness, on the other hand, refers to the capacity of a blockchain or smart contract language to simulate any Turing machine, which means it can resolve any computational problem given enough time and resources. Both those terms are used to represent smart contract.
Figure 2 illustrates the complete process of articles identifications. 504 conference and journal articles were found using the specified search string. Out of these, 41 articles are included based on the exclusion and inclusion criteria mentioned in table I. As for gray literature we focused on blog posts found within the first eight pages of Google search results. The search strings used included ”ethereum smart contract vulnerabilities and mitigation” and ”ethereum smart contract analysis tools”. Out of numerous blogs reviewed, eight posts were found to meet our established criteria. To assist the quality of the gray literature we have taken into consideration the following aspects:
@@
The expertise of the author in the field of smart contract security such as job position.
@@
- Studies related to vulnerabilities in Ethereum: Vulnerabilities explication and/or mitigation techniques.
- Studies related to the security improvement of Ethereum smart contracts. - Studies related to vulnerability detection tools in Ethereum smart contracts. - Studies that introduce open-source tools to detect the vulnerabilities of solidity smart-contracts. |
@@
- Results past the first eight Google pages, as we have noticed that after the eight pages the results are not relevant. - Data-sets, tweets, presentations. - Tools that use machine learning, as comparing machine learning tools need different criteria such as the used training models, accuracy, etc. . - Studies that are based on the improvement of blockchain infrastructure, rather than the smart contract programming - Study that is an older version of another paper that is previously been considered. |
III Vulnerabilities
@@
In this section, we present some of the common vulnerabilities of Ethereum smart contracts alongside real world attacks and prevention mechanisms.
III-A Reentrancy
III-A1 Description
@@
When implementing smart contracts, reentrancy is one of the most important vulnerability to address, also known as recursive call attack. A contract calling another contract, with external calls, will cause the stoppage of the calling contract’s execution and memory state until the called function returns a response. External calls are exposed in contract interfaces; hackers can use them to invoke a function within the contract numerous times, causing the contract to perform unanticipated tasks. This vulnerability occurs in a solidity smart contract performing critical tasks (e.g.,token transfer) before resolving the effects that should have been addressed (e.g., balance update).
III-A2 Implications on NFT Fractionalization
@@
Reentrancy vulnerabilities have already manifested significant impacts on decentralized finance (DeFi) protocols, illustrating potential risks for fractionalized NFT platforms. For instance, the dForce DeFi Protocol Hack [12] in February 2023, where an attacker exploited a reentrancy vulnerability in the Curve Finance vault on the Arbitrum and Optimism blockchains, part of the dForce protocol, led to the theft of approximately $3.6 million in assets. This attack underscores the danger reentrancy poses to smart contracts handling complex financial transactions, serving as a warning for platforms dealing with fractionalized NFTs. Furthermore, the CREAM Finance Hack [13] in August 2021 and the Siren Protocol Hack [14] in September 2021, with losses of $18.8 million and $3.5 million respectively, further exemplify the ongoing threat of reentrancy attacks to the blockchain and DeFi sectors.
@@
Reentrancy attacks can distort market dynamics by unfairly redistributing assets, leading to market manipulation and loss of liquidity. For fractionalized NFTs, where valuation depends on both the perceived value of the underlying asset and the platform’s integrity, such attacks can result in volatile price swings and diminished market confidence.
III-A3 Protection Measures
@@
It is recommended to use send() and transfer() methods instead of the general call() method to transfer money. They are deemed safer because they have a gas limit of 2,300 gas. Whereas, the method call() does not have a gas limit and forwards the remaining gas to the target address. The gas limit prohibits the target contract from making costly external function calls. The checks-effects-interactions pattern [15] is the most reliable approach to prevent reentrancy attacks. This pattern defines how the code of a function should be organized to minimize undesirable side effects and execution behaviors. The programmer needs to include all checks which usually consist of assert() and need() modifiers. If these checks pass, the function will then resolve all of the effects of the contract ( e.g. balance update). The function should only communicate with other contracts once all state changes have been updated. Even if an attacker performs a recursive call to the initial function, the user will not be able to exploit the state of the contract since external functions are called last.
@@
Another protection mechanism against reentrancy attacks, is the use of mutex. A mutex locks down the contract state. Mutex allows only one execution of a critical section of code at a time using lock mechanism. Only the lock’s owner has the capacity to modify it. When an attacker tries to perform a reentrancy attack, the state will be locked by the previous call and the balance can not be updated. Mutex must be treated with care to ensure that the lock can be released. ReentrancyGuard is OpenZeppelin’s own mutex implementation [16]. This library offers a nonReentrant modifier that secures an external function with a mutex.
III-A4 Example
@@
III-B Front Running
III-B1 Description
Front-running, also known as transaction ordering dependency, occurs when the execution logic depends on the order of the transactions submitted to it. The order of transactions in Ethereum is determined by the miner [17]. The transaction is visible to the network before being executed. The participants can exploit from this visibility by sending transactions with a higher gas price to be included first.
III-B2 Implications on NFT Fractionalization
@@
1. Manipulation of Share Prices: Malicious actors could exploit transaction ordering to manipulate the market for fractional tokens, buying up tokens at lower prices before a large transaction increases their value, or conversely, selling tokens to depress prices ahead of a significant sell order.
@@
III-B3 Protection Measures
The best solution to protect against front running vulnerability, is to remove the advantage of transaction ordering from the application. A solution is to remove the importance of time. Another possible solution is to use a commit-reveal hash scheme where the participant submits the hash of the answer instead of the answer. The contract then stores the hash and the address of the sender; the answer is revealed only after all all the responses are submitted.
III-B4 Example
Listing 5 illustrates a smart contract managing the sale of fractional tokens of an NFT, vulnerable to front-running. In this scenario, an attacker could observe pending purchase transactions and execute their own purchase with a higher gas fee, securing shares at the current price before a large purchase increases their value, thereby disadvantaging legitimate buyers.
To mitigate the vulnerability and ensure fair transaction processing, the contract can be revised to include a commit-reveal scheme. This modification requires buyers to commit to a purchase without revealing the exact quantity or price initially, followed by a reveal phase where the transaction details are disclosed. Mitigating the front-running attack as showing in Listing 6.
III-C Arithmetic
III-C1 Description
@@ III-C2 Implications on NFT Fractionalization
In the context of NFT fractionalization, arithmetic vulnerabilities pose unique challenges. For instance when distributing sales revenue among shareholders, overflow or underflow errors can lead to incorrect allocation of funds. This not only affects the fairness and accuracy of fees distribution but also exposes the platform to potential exploitation. Arithmetic vulnerabilities, particularly overflows and underflows, have had substantial consequences in the decentralized finance (DeFi) sector, highlighting the associated risks for platforms dealing with fractionalized NFTs. A prominent example includes the mintToken function in the Coinstar (CSTR) Ethereum token’s smart contract [19], which suffered from an integer overflow. This flaw permitted the contract owner to arbitrarily adjust the balance of any user to any chosen value. An other example is, the 4chan gang group experienced a substantial loss of $2.3 million due to an underflow vulnerability in the ERC-20 token implementation of PoWH (Proof of Weak Hands), which allowed an attacker to exploit this flaw for financial gain [20].
III-C3 Protection Measures

### v2

@@ Abstract
Ethereum smart contracts are highly powerful; they are immutable and retain massive amounts of tokens. However, smart contracts keep attracting attackers to benefit from smart contract flaws and Ethereum’s unexpected behavior. Thus, methodologies and tools have been proposed to help implement secure smart contracts and to evaluate the security of smart contracts already deployed. Most related surveys focus on tools without discussing the logic behind them; in addition, they assess the tools based on papers rather than testing the tools and collecting community feedback. Other surveys lack guidelines on how to use tools specific to smart contract functionalities. This paper presents a literature review combined with an experimental report that aims to assist developers in developing secure smarts, with a novel emphasis on the challenges and vulnerabilities introduced by NFT fractionalization by addressing the unique risks of dividing NFT ownership into tradeable units called fractions. It provides a list of frequent vulnerabilities and corresponding mitigation solutions. In addition, it evaluates the community’s most widely used tools by executing and testing them on sample smart contracts. Finally, a comprehensive guide on how to implement secure smart contracts is presented.
@@
Blockchain, Ethereum, Smart contracts, Formal verification, Semantic verification, Fuzzing, Software security, and Software quality.1 Introduction
@@
Blockchain technology has exploded in popularity over the years due to its immutability, security, and transparency in a permissionless and decentralized environment. One of the well-known and used implementations is Ethereum. As of 31 May 2023, the Ethereum cryptocurrency’s market capital surpassed $220 billion, with millions of transactions being executed every day [1]. Ethereum uses Turing Complete blockchain technology [2], allowing developers to implement smart contracts. Smart contracts are executable programs stored in the blockchain written primarily in solidity [3]. Using smart contracts facilitates the execution of pre-defined terms without consulting third parties in an anonymous, transparent, and tempered manner.
@@
The Ethereum blockchain is the foundation for numerous financial projects. Examples include (a) platforms for decentralized finance (DeFi), which enable users to access various financial services like lending and borrowing, trading, and insurance without intermediates. (b) Stablecoins, cryptocurrencies whose value is tied to that of a fiat currency or other asset, like gold or the US dollar; (c) Security token offerings (STOs), which are virtual assets that simulate ownership of a physical asset like a stock or piece of real estate. (d) Non-fungible tokens (NFTs), which are unique digital assets that might signify ownership of a physical or digital object, like a work of art or a collectible, (e) Amidst this variety, NFT fractionalization emerges as a noteworthy trend, dividing ownership of NFTs into smaller, more accessible units, thus broadening participation in the digital asset market. This value attracts attackers to exploit different vulnerabilities related to implementing smart contracts to steal cryptocurrencies or tamper with assets. For instance, a recent attack was on May 2023 where Level Finance Exchange announced the loss of more than 214,000 $LVL tokens, approximated to be $1.01 million [4], caused by an attack on its smart contract by manipulating a recursive calling vulnerability. There are various vulnerabilities and causes, such as arithmetic overflows, reentrancy, inadequate randomness, calling an unknown third-party contract code, and failing to check the return status of external calls. The runtime environment that supports contract code execution provides more attack points. One example is when malevolent miners pick and choose which transactions are included in a mined block or in what order they are included.
@@
Since vulnerabilities in smart contracts can result in significant financial loss, multiple tools are proposed to identify vulnerabilities in Solidity smart contracts. Detection methods include (a) Symbolic execution when the program is abstractly executed to cover various possible inputs; (b) Fuzzy testing by injecting several invalid inputs; (c)Taint analysis by checking an input flow; and (d) Formal verification, when the behavior of the smart contract is checked mathematically using a formal model.
@@
In this paper, we conduct a literature review and an experimental report. We present the most common vulnerabilities in Solidity smart contracts, including those posed by the fractionalization of NFTs and the corresponding mitigation schemes. We also overview and compare the most popular schemes and tools used to detect smart contract vulnerabilities. Finally, we propose a set of guidelines for auditing smart contracts.
@@
In this survey, we searched 2 popular databases, Engineering Village and Scopus, and web articles written by the Ethereum community Concerning vulnerabilities in Solidity smart contracts, mitigation tools, and guidelines. Moreover, we experimentally studied 5 tools widely used to detect vulnerabilities in smart contracts, namely Oyente [5], Slither [6], Mythril [7], Manticore [8] and Echidna [9].
@@
The rest of this paper is organized as follows: Section 2 outlines the methodology we used to produce this survey. Section 3 presents the most common vulnerabilities of smart contracts and the corresponding mitigation schemes. Section 4 presents standard methodologies for detecting smart contract vulnerabilities. Section 5 introduces the most valuable tools to detect vulnerabilities in smart contracts. Section 6 presents a set of guidelines for auditing smart contracts. Section 7 presents an overview of existing related surveys. Finally, Section 8 concludes the paper.
2 Survey Methodology
@@
RQ3: What are the existing methodologies used to detect the vulnerabilities of smart contracts, and how do they compare?
@@
To address these questions, we conducted a Multivocal Literature Review (MLR)[10] for data preparation; This technique incorporates both gray literature and white literature. Gray literature includes blogs, videos, and forums; it is usually written by practitioners from industry and academia. It’s not peer-reviewed. White literature contains peer-reviewed research articles from journals. We chose the published studies in journals and conferences with high-impact factors and competitive acceptance rates. We also check the citation count of the studies being chosen on Google Scholar to evaluate their impact on the evolution of this emerging paradigm. The rationale behind choosing MLR is that the field of smart contracts security is still relatively new; thus, including literature from practitioners helps provide a complete overview. Moreover, the feedback provided by smart contract developers is crucial in selecting the most useful tools.
@@
We defined a search string presented in Listing 1 to collect the following resources: (a) Review articles about smart contract security; (b) Papers that discuss vulnerabilities in Ethereum smart contracts; (c) Papers that cover detection methodologies or mitigation schemes of smart contract vulnerabilities; and (d) Papers about tools that detect vulnerabilities in smart contacts. We have only included papers containing the keywords in the subject, title, or abstract. The term ”opcode” refers to the basic instructions carried out by the blockchain’s virtual machine, which regulates the logic and operations of the smart contract. Turing completeness, on the other hand, refers to the capacity of a blockchain or smart contract language to simulate any Turing machine. This means it can resolve any computational problem given enough time and resources. Both those terms are used to represent smart contracts.
Figure 2 illustrates the complete process of article identifications. 504 conference and journal articles were found using the specified search string. 41 articles are included based on the exclusion and inclusion criteria mentioned in table I. As for gray literature, we focused on blog posts found within the first eight pages of Google search results. The search strings include ”Ethereum smart contract vulnerabilities and mitigation” and ”Ethereum smart contract analysis tools.” Out of the numerous blogs reviewed, eight posts met our established criteria. To assist the quality of the gray literature, we have taken into consideration the following aspects:
@@
The author’s expertise in smart contract security, such as job position.
@@
- Studies related to vulnerabilities in Ethereum: Vulnerabilities explication and mitigation techniques.
@@
3 Vulnerabilities
@@
In this section, we present some of the common vulnerabilities of Ethereum smart contracts alongside real-world attacks and prevention mechanisms.
3.1 Reentrancy
3.1.1 Description
@@
Reentrancy is one of the most critical vulnerabilities to address when implementing smart contracts, also known as a recursive call attack. A contract calling another contract, with external calls, will cause the stoppage of the calling contract’s execution and memory state until the called function returns a response. External calls are exposed in contract interfaces; hackers can use them to invoke a function within the contract numerous times, causing the contract to perform unanticipated tasks. This vulnerability occurs in a solidity smart contract performing critical tasks (e.g., token transfer) before resolving the effects that should have been addressed (e.g., balance update).
3.1.2 Implications on NFT Fractionalization
@@
Reentrancy vulnerabilities have significantly impacted decentralized finance (DeFi) protocols, illustrating potential risks for fractionalized NFT platforms. For instance, the dForce DeFi Protocol Hack [12] in February 2023, where an attacker exploited a reentrancy vulnerability in the Curve Finance vault on the Arbitrum and Optimism blockchains, part of the dForce protocol, led to the theft of approximately $3.6 million in assets. This attack underscores the danger reentrancy poses to smart contracts handling complex financial transactions, serving as a warning for platforms dealing with fractionalized NFTs. Furthermore, the CREAM Finance Hack [13] in August 2021 and the Siren Protocol Hack [14] in September 2021, with losses of $18.8 million and $3.5 million, respectively, further exemplify the ongoing threat of reentrancy attacks to the blockchain and DeFi sectors.
@@
Reentrancy attacks can distort market dynamics by unfairly redistributing assets, leading to market manipulation and loss of liquidity. For fractionalized NFTs, where valuation depends on the underlying asset’s perceived value and the platform’s integrity, such attacks can result in volatile price swings and diminished market confidence.
3.1.3 Protection Measures
@@
It is recommended to use send() and transfer() methods instead of the general call() method to transfer money. They are deemed safer because they have a gas limit of 2,300 gas. At the same time, the method call() does not have a gas limit and forwards the remaining gas to the target address. The gas limit prohibits the target contract from making costly external function calls. The checks-effects-interactions pattern [15] is the most reliable approach to prevent reentrancy attacks. This pattern defines how the code of a function should be organized to minimize undesirable side effects and execution behaviors. The programmer needs to include all checks, which usually consist of assert() and need() modifiers. If these checks pass, the function will resolve all of the effects of the contract ( e.g., balance update). The function should only communicate with other contracts once all state changes have been updated. Even if an attacker performs a recursive call to the initial function, the user cannot exploit the state of the contract since external functions are called last.
@@
Another protection mechanism against reentrancy attacks is the use of mutex. A mutex locks down the contract state. Mutex allows only one execution of a critical code section at a time using a lock mechanism. Only the lock’s owner can modify it. When an attacker tries to perform a reentrancy attack, the previous call will lock the state, and the balance can not be updated. Mutex must be treated with care to ensure the lock can be released. ReentrancyGuard is OpenZeppelin’s own mutex implementation [16]. This library offers a nonReentrant modifier that secures an external function with a mutex.
3.1.4 Example
@@
3.2 Front Running
3.2.1 Description
Front-running, also known as transaction ordering dependency, occurs when the execution logic depends on the order of the submitted transactions. The miner [17] determines the order of transactions in Ethereum. The transaction is visible to the network before being executed. The participants can exploit this visibility by sending transactions with a higher gas price to be included first.
3.2.2 Implications on NFT Fractionalization
@@
1. Manipulation of Share Prices: Malicious actors could exploit transaction ordering to manipulate the market for fractional tokens, buying up tokens at lower prices before a significant transaction increases their value or, conversely, selling tokens to depress prices ahead of a significant sell order.
@@
3.2.3 Protection Measures
The best solution to protect against front-running vulnerability is to remove the advantage of transaction ordering from the application. A solution is to remove the importance of time. Another possible solution is to use a commit-reveal hash scheme where the participant submits the hash of the answer instead of the answer. The contract then stores the hash and the sender’s address; the answer is revealed only after all the responses are submitted.
3.2.4 Example
Listing 5 illustrates a smart contract managing the sale of fractional tokens of an NFT, vulnerable to front-running. In this scenario, an attacker could observe pending purchase transactions and execute their purchase with a higher gas fee, securing shares at the current price before a large purchase increases their value, thereby disadvantaging legitimate buyers.
The contract can be revised to include a commit-reveal scheme to mitigate the vulnerability and ensure fair transaction processing. This modification requires buyers to commit to a purchase without initially revealing the exact quantity or price. This is followed by a reveal phase where the transaction details are disclosed as shown in Listing 6.
3.3 Arithmetic
3.3.1 Description
@@ 3.3.2 Implications on NFT Fractionalization
In the context of NFT fractionalization, arithmetic vulnerabilities pose unique challenges. For instance, when distributing sales revenue among shareholders, overflow or underflow errors can lead to incorrect allocation of funds. This affects the fairness and accuracy of fee distribution and exposes the platform to potential exploitation. Arithmetic vulnerabilities, particularly overflows and underflows, have had substantial consequences in the decentralized finance (DeFi) sector, highlighting the associated risks for platforms dealing with fractionalized NFTs. A prominent example includes the mintToken function in the Coinstar (CSTR) Ethereum token’s smart contract [19], which suffered from an integer overflow. This flaw permitted the contract owner to adjust any user’s balance to any chosen value arbitrarily. Another example is the 4chan gang group experienced a substantial loss of $2.3 million due to an underflow vulnerability in the ERC-20 token implementation of PoWH (Proof of Weak Hands), which allowed an attacker to exploit this flaw for financial gain [20].
3.3.3 Protection Measures